OpenStudy (anonymous):

7(x - 2) + 3(x - 2) = 5(x - 6) in fraction

OpenStudy (anonymous):

find for x

OpenStudy (lilg132):

7x - 14 + 3x - 6 = 5x - 30 7x + 3x - 5x = 14 + 6 - 30 5x = -10 x = -2

OpenStudy (anonymous):

that cant be the right answer

OpenStudy (anonymous):

the answer has to be a fraction

OpenStudy (lilg132):

7(x - 2) + 3(x - 2) = 5(x - 6) 7x - 14 + 3x - 6 = 5x - 30 7x + 3x - 14 - 6 = 5x - 30 10x - 20 = 5x - 30 30 - 20 = 5x - 10x 10 = 5x 2 = x no it cant be a fraction answer

OpenStudy (anonymous):

it has to!

OpenStudy (anonymous):

can you put 7.6 into a fraction?

OpenStudy (lilg132):

\[\frac{38}{5}\]

OpenStudy (anonymous):

okay the answer was -4\[\left\{ 2 \right\}\left\{ 5 \right\}\]

OpenStudy (anonymous):

2/5

OpenStudy (lilg132):

how did you get that

OpenStudy (anonymous):

i didnt i got it wrong and they showed me the answer.. :/

OpenStudy (lilg132):

lol so you dont know how it was worked out?

OpenStudy (anonymous):

Nope

OpenStudy (lilg132):

you should find out :) for your own knowledge

OpenStudy (anonymous):

lol yeah i know but don't worry i have 4 more.. :P